Calculate the energy, in joules, of a C5+ ion with an electron in the n=3 state.
Answer #1

Answer) Energy of C5+ ion with electron in n= 3 state is -8.72 * 10-18 J

C⁵⁺ ion is an atom with a single electron. So the predictions of the Bohr-model hold for this ion. According to Bohr model the energy of an electron orbiting at level n around a nucleus with Z protons is given by:
En= - RH · Z² / n²

En is energy in nth state

RH is Rydberg constant = 2.18 * 10-18J

Z is atomic number

Z of carbon = 6

n = 3 ( given)

Putting values in formula we get ,

E3 = - [(2.18 * 10-18 J ) * (6)2 ]/(3)2

E3 = - [(2.18 * 10-18 J ) * (36)]/9

E3 = - (2.18 * 10-18 J ) * (4)

E3 = - 8.72 * 10-18 J